OpenStudy (anonymous):

want some medals?.. log2(2m+4)−log2(m−1)=3

OpenStudy (anonymous):

can i redeem them for valuable gifts?

OpenStudy (anonymous):

i want some

OpenStudy (anonymous):

\[\log_{2}(2m+4) - \log_{2}(m-1)=3\]

OpenStudy (anonymous):

this time it is a subtraction so write \[\log_2(\frac{2m+4}{m-1})\]=3\] then solve \[\frac{2m+4}{m-1}=2^3\]

OpenStudy (anonymous):

thats what it was

OpenStudy (anonymous):

can i redeem them for valuable gifts?

OpenStudy (anonymous):

lol@satllite : )

OpenStudy (anonymous):

\[\log_{2}(2m+4)-\log_{2}(m-1)= 3\] \[\frac{2m+4}{m-1}=8\] 2m+4=8m-8 6m=12 m=12/6=2